how many grams of pure salt must be added to 40g of a 20% salt solution to make a solution that is 36% salt

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

how many grams of pure salt must be added to 40g of a 20% salt solution to make a solution that is 36% salt
-------------------------------
Equation:
salt + salt = salt
0.20*40 + 1.00x = 0.36*(40+x)
Multiply thru by 100 to get:
20*40 + 100x = 36*40 + 36x
64x = 16*40
x = (1/4)*40
x = 10 grams (amt. of pure salt needed for the mixture)
==================
Cheers,
Stan H.

Answer by mananth(16946)

how many grams of pure salt must be added to 40g of a 20% salt solution to make a solution that is 36% salt
..
20% -----40g
100% -----x gm
Mixture ---36%-----(x+40)gm
..
0.2*40+x=0.36*(x+36)
8+x=0.36x+12.96
x-0.36x=12.96-8
0.64x= 4.96
x=4.96/0.64
= 7.75 gm of pure salt.
...
m.ananth@hotmail.ca